Jenny does ironing for a dry cleaner. She is paid $150.00 per month plus .25 per shirt and $1.50 per pair of slacks. Last month she ironed a total of 300 pieces and she made $325.00. How many shirts did she irom? How many slacks did she iron?

Answer by mananth(16946)

shirts ----------x
slacks------------y
x+y=300...........................(1)
0.25x+1.50y+150=325
multiply by 100
25x+150y+15000=32500
25x+150y=17500....................(2)

x= shirts
y= slacks

1 x + 1 y = 300 .............1
Total value
25 x + 150 y = 17,500 .............2
Eliminate y
multiply (1)by -150
Multiply (2) by 1
-150 x -150 y = -45,000
25 x + 150 y = 17,500
Add the two equations
-125 x = -27,500
/ -125
x = 220
plug value of x in (1)
1 x + 1 y = 300
220 + y = 300
y = 300 -220
y = 80
y = 80
x= 220 shirts
y= 80 slacks
m.ananth@hotmail.ca